Use a normal approximation to find the probability of the indicated number of voters. In this​ case, assume that 137 eligible voters aged​ 18-24 are randomly selected. Suppose a previous study showed that among eligible voters aged​ 18-24, 22% of them voted.
Probability that fewer than 32 voted
The probability that fewer than 32 of 137 eligible voters voted is __.
(Round to four decimal places as​ needed.)

Answer by Boreal(15235)

np and n(1-p) are both >10
mean is np=137*0.22=30.14
variance is 30.14*(0.78)=23.5092
sd is sqrt(V)=4.85
z < (31.5-30.14)/4.85, using the continuity correction factor.
< 1.36/4.85 or 0.28
That probability is 0.6103